# MoORE: SVD-based Model MoE-ization for Conflict- and Oblivion-Resistant Multi-Task Adaptation

### v4

@@ 3.2 Comparisons with Existing MoE-based Multi-Task Adaptation Methods
LoRAMoE [14] MDMD $M(D_{\text{out}}+D)r$ $\mathcal{O}(D_{\text{out}}D+M(D+D_{\text{out}}r+Dr))$
MixLoRA [30] MDMD $M(D_{\text{out}}+D)r$ $\mathcal{O}(D_{\text{out}}D+MD+(D_{\text{out}}r+Dr))$
MoSLD [81] MDMD $(D_{\text{out}}+MD)r$ $\mathcal{O}(D_{\text{out}}D+D_{\text{out}}r+M(Dr+D))$
HydraLoRA [60] MDMD $(MD_{\text{out}}+D)r$ $\mathcal{O}(D_{\text{out}}D+Dr+M(D_{\text{out}}r+D))$
MTL-LoRA [73] MKMK $(MD_{\text{out}}+Kr+D)r$ $\mathcal{O}(D_{\text{out}}D+Dr+MD_{\text{out}}r+r^{2})$
OMoE [18] MDMD $M(D_{\text{out}}+D)r$ $\mathcal{O}(D_{\text{out}}(D+M^{2})+M(D+D_{\text{out}}r+Dr))$
@@ 3.3 Comparisons with Existing SVD-based Fine-Tuning Methods
SVF [56] $\bm{U}\text{diag}(\sigma)\bm{V}^{\top}$ Nonparametric $\sigma$ Original architecture: $\bm{W}_{\text{new}}$
SVDiff [21] $\bm{U}\text{diag}(\text{ReLU}(\sigma+\delta))\bm{V}^{\top}$ Nonparametric $\delta$ Original architecture: $\bm{W}_{\text{new}}$
SVFT [33] $\bm{U}(\text{diag}(\sigma)+\bm{M})\bm{V}^{\top}$ Nonparametric $\bm{M}$ Original architecture: $\bm{W}_{\text{new}}$
KaSA [64] $\bm{W}_{lr}+\Delta\bm{U}\text{diag}(\Delta\sigma)\Delta\bm{V}^{\top}$ Nonparametric $\Delta\bm{U},\Delta\sigma,\Delta V^{\top}$ Original architecture: $\bm{W}_{\text{new}}$
SVFCL [69] $\bm{U}\text{diag}(\sigma+\sum_{i=0}^{t}\sigma_{i})\bm{V}^{\top}$ Nonparametric $\sigma_{i}$’s Original architecture: $\bm{W}_{\text{new}}$
@@
SVF [56] No No No
SVDiff [21] No No No
SVFT [33] No No No
KaSA [64] No No No
SVFCL [69] Yes (incremental learning) No Partially discussed
@@
Our method employs SVD to perform model MoE-ization, constructing a complete MoE architecture. This fundamentally differs from existing SVD-based fine-tuning methods [48, 56, 21, 33, 64, 69]. To more clearly illustrate the differences between MoORE and these methods, we select five recent ones for comparison, including SVF [56], SVDiff [21], SVFT [33], KaSA [64], and SVFCL [69]. Tables 3 and 4 compare different methods in terms of implementation and application, respectively.
@@ 4 Experiments
To demonstrate the effectiveness of MoORE in multi-task adaptation, we apply three MTL datasets and conduct comprehensive experiments on them. Representative results are shown in Figure 4(a) and the following content. More experimental details, e.g., the basic information of datasets, hyperparameter settings, ablation studies, routing weight analysis, and numerical results associated with figures, are shown in the Appendix.
@@ 4.1 Implementation Details
Base model and baselines. In the following experiments, we utilize LLaMA-3.1 8B [^2] https://huggingface.co/meta-llama/Llama-3.1-8B-Instruct [20] as the base model and adapt it by various multi-task adaptation methods. In particular, we compare MoORE with LoRA [25] and the methods incorporating low-rank adapters as MoEs, including LoRAMoE [14], MoSLD [81], MTL-LoRA [73], HydraLoRA [60], and MixLoRA [30]. We implement the MoE architectures of the baselines mainly based on their default settings. For a fair comparison, we modify some baselines’ hyperparameters to make the number of learnable parameters comparable for each method. For MoORE, we MoE-ize all linear layers of LLaMA-3.1 8B, including the “QKVOQKVO” modules of attention layers and the weight matrices of FFNs.
Two datasets for evaluating conflict-resistance. We consider two MTL datasets for commonsense reasoning (CSR) and natural language understanding (NLU), respectively. The CSR-MTL dataset is constructed by nine tasks, including ARC-Challenge (ARC-C), ARC-Easy (ARC-E) [9], OpenBookQA (OBQA) [45], PIQA [5], SocialIQA (SIQA) [54], BoolQ [8], Hellaswag (HellaS) [77], Winogrande (WinoG) [53], and CommonsenseQA (CSQA) [59]. These tasks are widely used to evaluate LLMs on various commonsense reasoning challenges, ranging from genuine grade-school level science questions to physical commonsense reasoning. The NLU-MTL dataset consists of seven tasks from GLUE [63], including CoLA, SST-2, MRPC, QQP, MNLI, QNLI, and RTE. These tasks are applied to evaluate the natural language understanding capabilities of LLMs, including natural language inference, textual entailment, sentiment analysis, semantic similarity, and so on.
One dataset for evaluating oblivion-resistance. In addition, we construct one more dataset, called OR-MTL, for evaluating the oblivion-resistance of different methods. The dataset includes seven tasks, including MMLU [24, 23], IFEval [83], BIG-Bench Hard (BBH) [58], GPQA [51], HumanEval [7], MBPP [4], and GSM-8K [10]. The base model, LLaMA-3.1 8B, achieves encouraging performance in these tasks. After adapting it on CSR-MTL, we record the performance of the adapted models in OR-MTL and assess the ability of different methods to mitigate task oblivion.
Experimental settings. When adapting the pre-trained model on CSR-MTL and NLU-MTL, we set the training epoch to $2$ and $5$, respectively. The learning rate is set to $3\times 10^{-4}$, with AdamW [39] as the optimizer. For CSR-MTL, we set the batch size to $8$, whereas for NLU-MTL, we set the batch size to $64$. Both training and testing are conducted on one NVIDIA A100 GPU.
@@
LoRA [25] 2.09% 77.56 85.77 70.43 81.60 82.97 76.00 93.00 71.11 77.40 79.54
MixLoRA [30] 3.00% 79.18 87.50 72.02 86.60 87.38 78.86 93.65 77.35 80.43 82.55
MoSLD [81] 1.49% 80.29 86.87 74.16 86.40 88.58 79.73 95.03 80.58 80.34 83.55
HydraLoRA [60] 2.72% 79.27 88.09 74.34 85.60 88.96 79.84 95.36 83.03 80.10 83.84
MTL-LoRA [73] 2.69% 80.55 88.38 75.26 85.80 88.41 80.45 95.57 81.61 80.75 84.09
LoRAMoE [14] 2.19% 80.97 88.51 74.37 86.20 89.45 80.91 95.26 81.29 82.06 84.34
@@
LoRA [25] 2.09% 39.69 86.20 80.88 88.16 87.80 80.14 90.60 79.07
MixLoRA [30] 3.00% 61.71 89.98 82.84 94.47 90.11 85.56 95.53 85.74
MoSLD [81] 1.49% 58.55 90.34 85.29 94.95 90.08 89.53 96.33 86.44
MTL-LoRA [73] 2.69% 60.65 90.26 87.01 95.26 91.95 90.98 96.10 87.46
HydraLoRA [60] 2.72% 64.92 89.97 86.76 95.31 91.62 87.73 96.44 87.54
LoRAMoE [14] 2.19% 62.84 91.01 88.97 95.55 91.01 90.61 96.56 88.08
@@ 4.2 Performance in Conflict-Resistance
Conflict-resistance regarding task number and difficulty. To further compare and analyze the conflict-resistance capabilities of different methods, we conduct comparative experiments on CSR-MTL by varying the number and difficulty of the tasks. In particular, for each task in CSR-MTL, we first calculate the average of all the methods’ performance based on the results in Table 5. The average performance of the methods in a task measures the difficulty of the task for the base model — the lower the average performance is, the more difficult the task is. Then, we sort the tasks in ascending order based on their difficulty. Finally, we adapt the base model for the top-$K$ tasks, K=2,,9K=2,...,9, and show the performance of different methods in Figure LABEL:fig:increment1. With the increase of task number and difficulty, all the methods suffer performance degradation because $i)$ task conflict becomes severe as the number of tasks increases, and ii)ii) difficult tasks are more likely to have conflicts with other tasks. Notably, MoORE outperforms all other baselines across all settings.
(a) The loss of performance in OR-MTL. (a-g) The gray dashed line represents the performance of LLaMA-3.1 8B before adaptation. (h) The overall performance degradation across all tasks.
@@ 4.3 Performance in Oblivion-Resistance
To demonstrate the superiority of MoORE in mitigating task oblivion, we compare various methods in the Language Model Evaluation Harness framework [19]. In particular, we first adapt the base model on CSR-MTL using different methods. Then, we evaluate the adapted models on OR-MTL, comparing them with the base model. Figure 12(a) shows that MoORE consistently mitigates task oblivion across all seven tasks of OR-MTL, with an average performance drop of only 1.31% compared to the base model. It significantly outperforms the other adaptation methods. On HumanEval, MoORE achieves performance exceeding the original model, with an improvement of 9.75%. Similarly, LoRAMoE and MTL-LoRA also obtain slight improvements of 4.27% and 0.16%, respectively. This intriguing phenomenon may imply that the datasets in CSR-MTL have some relevance to HumanEval, providing information and capabilities beneficial for solving this task.
Oblivion-resistance regarding task number and difficulty. Furthermore, we conduct experiments to investigate how the ability of oblivion-resistance changes with increased task number and difficulty. The results in Figure LABEL:fig:increment2 show that after adapting to the top-3 tasks of CSR-MTL, MoORE and its strongest competitor, LoRAMoE [14], achieve comparable performance on OR-MTL. However, with the increase of task number and difficulty, the performance of LoRAMoE changes slightly while the performance of MoORE increases and becomes superior to that of LoRAMoE. This result demonstrates that MoORE has stronger oblivion-resistance than its competitors.
@@
The generalization to unseen task IDs. The computation of task-level weights requires task IDs as input, but the ID of a new task is unavailable during inference. Therefore, we use the average of the existing task ID embeddings as the embedding of the new task. Specifically, after adapting the base model on CSR-MTL, MoORE only learns the task IDs of the nine tasks in CSR-MTL. During evaluation on OR-MTL, we use the average of the nine task ID embeddings as the task ID embedding of each new task. The results in Figure LABEL:fig:increment2 and 12(a) validate the effectiveness of this approach.
In addition, this experiment also explains the result in Figure LABEL:fig:increment2. In particular, the more tasks considered in the adaptation phase, the more likely some tasks correlate with those covered in the pre-training phase. As a result, increasing the number of tasks in the adaptation phase helps enhance the oblivion-resistance of MoORE.
@@ 4.4 Inference Efficiency
To compare inference efficiency, we adapt the base model using MoORE and other baselines, and evaluate their inference time on GPQA. For a fair comparison, all experiments are conducted on a single NVIDIA A100 GPU with the same batch size. The results in Figure LABEL:fig:speed show the average inference time per batch for each method, which indicates that MoORE achieves superior inference efficiency compared to most baselines and is comparable to MixLoRA. Compared to the base model (i.e., the gray dashed line), MoORE moderately increases inference time.
@@ 5 Conclusion and Future Work
Limitations and future work. As shown in Figure LABEL:fig:speed, the inference efficiency of MoORE is comparable to that of baselines. However, further reducing its complexity for large-scale applications is challenging. In particular, MoORE applies many more experts than classic MoE models, and the analytic experiments in Appendix C.4 show that it learns dense weights for the experts. In addition, besides imposing orthogonal adapters, we need to find a more effective solution to enhance model capacity in broader scenarios. In theory, the results in Figures LABEL:fig:increment1 and LABEL:fig:increment2 show that increasing the number of tasks leads to a higher risk of task conflict while helping mitigate task oblivion. How to understand this “trade-off”? How can the limitation in multi-task adaptation be quantified regarding the number of tasks? These are still open problems. Based on the above analysis, we plan to explore efficient implementations of MoORE and study its theoretical properties in the future.

### v5

@@ 3.2 Comparisons with Existing MoE-based Multi-Task Adaptation Methods
LoRAMoE dou2024loramoe MDMD $M(D_{\text{out}}+D)r$ $\mathcal{O}(D_{\text{out}}D+M(D+D_{\text{out}}r+Dr))$
MixLoRA li2024mixlora MDMD $M(D_{\text{out}}+D)r$ $\mathcal{O}(D_{\text{out}}D+MD+(D_{\text{out}}r+Dr))$
MoSLD zhao2025mosld MDMD $(D_{\text{out}}+MD)r$ $\mathcal{O}(D_{\text{out}}D+D_{\text{out}}r+M(Dr+D))$
HydraLoRA tian2024hydralora MDMD $(MD_{\text{out}}+D)r$ $\mathcal{O}(D_{\text{out}}D+Dr+M(D_{\text{out}}r+D))$
MTL-LoRA yang2025mtl MKMK $(MD_{\text{out}}+Kr+D)r$ $\mathcal{O}(D_{\text{out}}D+Dr+MD_{\text{out}}r+r^{2})$
OMoE feng2025omoe MDMD $M(D_{\text{out}}+D)r$ $\mathcal{O}(D_{\text{out}}(D+M^{2})+M(D+D_{\text{out}}r+Dr))$
@@ 3.3 Comparisons with Existing SVD-based Fine-Tuning Methods
SVF sun2022singular $\bm{U}\text{diag}(\sigma)\bm{V}^{\top}$ Nonparametric $\sigma$ Original architecture: $\bm{W}_{\text{new}}$
SVDiff han2023svdiff $\bm{U}\text{diag}(\text{ReLU}(\sigma+\delta))\bm{V}^{\top}$ Nonparametric $\delta$ Original architecture: $\bm{W}_{\text{new}}$
SVFT lingam2024svft $\bm{U}(\text{diag}(\sigma)+\bm{M})\bm{V}^{\top}$ Nonparametric $\bm{M}$ Original architecture: $\bm{W}_{\text{new}}$
KaSA wang2024kasa $\bm{W}_{lr}+\Delta\bm{U}\text{diag}(\Delta\sigma)\Delta\bm{V}^{\top}$ Nonparametric $\Delta\bm{U},\Delta\sigma,\Delta V^{\top}$ Original architecture: $\bm{W}_{\text{new}}$
SVFCL wang2025singular $\bm{U}\text{diag}(\sigma+\sum_{i=0}^{t}\sigma_{i})\bm{V}^{\top}$ Nonparametric $\sigma_{i}$’s Original architecture: $\bm{W}_{\text{new}}$
@@
SVF sun2022singular No No No
SVDiff han2023svdiff No No No
SVFT lingam2024svft No No No
KaSA wang2024kasa No No No
SVFCL wang2025singular Yes (incremental learning) No Partially discussed
@@
Our method employs SVD to perform model MoE-ization, constructing a complete MoE architecture. This fundamentally differs from existing SVD-based fine-tuning methods peng2024sam ; sun2022singular ; han2023svdiff ; lingam2024svft ; wang2024kasa ; wang2025singular . To more clearly illustrate the differences between MoORE and these methods, we select five recent ones for comparison, including SVF sun2022singular , SVDiff han2023svdiff , SVFT lingam2024svft , KaSA wang2024kasa , and SVFCL wang2025singular . Tables 3 and 4 compare different methods in terms of implementation and application, respectively.
@@ 4 Experiments
To demonstrate the effectiveness of MoORE in multi-task adaptation, we apply three MTL datasets and conduct comprehensive experiments on them. Representative results are shown in Figure 1 and the following content. More experimental details, e.g., the basic information of datasets, hyperparameter settings, ablation studies, routing weight analysis, and numerical results associated with figures, are shown in the Appendix.
@@ 4.1 Implementation Details
Base model and baselines. In the following experiments, we utilize LLaMA-3.1 8B [^2] https://huggingface.co/meta-llama/Llama-3.1-8B-Instruct grattafiori2024llama3herdmodels as the base model and adapt it by various multi-task adaptation methods. In particular, we compare MoORE with LoRA hu2021lora and the methods incorporating low-rank adapters as MoEs, including LoRAMoE dou2024loramoe , MoSLD zhao2025mosld , MTL-LoRA yang2025mtl , HydraLoRA tian2024hydralora , and MixLoRA li2024mixlora . We implement the MoE architectures of the baselines mainly based on their default settings. For a fair comparison, we modify some baselines’ hyperparameters to make the number of learnable parameters comparable for each method. For MoORE, we MoE-ize all linear layers of LLaMA-3.1 8B, including the “QKVOQKVO” modules of attention layers and the weight matrices of FFNs.
Two datasets for evaluating conflict-resistance. We consider two MTL datasets for commonsense reasoning (CSR) and natural language understanding (NLU), respectively. The CSR-MTL dataset is constructed by nine tasks, including ARC-Challenge (ARC-C), ARC-Easy (ARC-E) allenai:arc , OpenBookQA (OBQA) OpenBookQA2018 , PIQA Bisk2020 , SocialIQA (SIQA) sap2019socialiqa , BoolQ clark2019boolq , Hellaswag (HellaS) zellers2019hellaswag , Winogrande (WinoG) sakaguchi2021winogrande , and CommonsenseQA (CSQA) talmor-etal-2019-commonsenseqa . These tasks are widely used to evaluate LLMs on various commonsense reasoning challenges, ranging from genuine grade-school level science questions to physical commonsense reasoning. The NLU-MTL dataset consists of seven tasks from GLUE wang2019glue , including CoLA, SST-2, MRPC, QQP, MNLI, QNLI, and RTE. These tasks are applied to evaluate the natural language understanding capabilities of LLMs, including natural language inference, textual entailment, sentiment analysis, semantic similarity, and so on.
One dataset for evaluating oblivion-resistance. In addition, we construct one more dataset, called OR-MTL, for evaluating the oblivion-resistance of different methods. The dataset includes seven tasks, including MMLU hendryckstest2021 ; hendrycks2021ethics , IFEval zhou2023instructionfollowingevaluationlargelanguage , BIG-Bench Hard (BBH) suzgun2022challenging , GPQA rein2024gpqa , HumanEval chen2021evaluating , MBPP austin2021program , and GSM-8K cobbe2021gsm8k . The base model, LLaMA-3.1 8B, achieves encouraging performance in these tasks. After adapting it on CSR-MTL, we record the performance of the adapted models in OR-MTL and assess the ability of different methods to mitigate task oblivion.
Experimental settings. When adapting the pre-trained model on CSR-MTL and NLU-MTL, we set the training epoch to $2$ and $5$, respectively. The learning rate is set to $3\times 10^{-4}$, with AdamW loshchilov2017decoupled as the optimizer. For CSR-MTL, we set the batch size to $8$, whereas for NLU-MTL, we set the batch size to $64$. Both training and testing are conducted on one NVIDIA A100 GPU.
@@
LoRA hu2021lora 2.09% 77.56 85.77 70.43 81.60 82.97 76.00 93.00 71.11 77.40 79.54
MixLoRA li2024mixlora 3.00% 79.18 87.50 72.02 86.60 87.38 78.86 93.65 77.35 80.43 82.55
MoSLD zhao2025mosld 1.49% 80.29 86.87 74.16 86.40 88.58 79.73 95.03 80.58 80.34 83.55
HydraLoRA tian2024hydralora 2.72% 79.27 88.09 74.34 85.60 88.96 79.84 95.36 83.03 80.10 83.84
MTL-LoRA yang2025mtl 2.69% 80.55 88.38 75.26 85.80 88.41 80.45 95.57 81.61 80.75 84.09
LoRAMoE dou2024loramoe 2.19% 80.97 88.51 74.37 86.20 89.45 80.91 95.26 81.29 82.06 84.34
@@
LoRA hu2021lora 2.09% 39.69 86.20 80.88 88.16 87.80 80.14 90.60 79.07
MixLoRA li2024mixlora 3.00% 61.71 89.98 82.84 94.47 90.11 85.56 95.53 85.74
MoSLD zhao2025mosld 1.49% 58.55 90.34 85.29 94.95 90.08 89.53 96.33 86.44
MTL-LoRA yang2025mtl 2.69% 60.65 90.26 87.01 95.26 91.95 90.98 96.10 87.46
HydraLoRA tian2024hydralora 2.72% 64.92 89.97 86.76 95.31 91.62 87.73 96.44 87.54
LoRAMoE dou2024loramoe 2.19% 62.84 91.01 88.97 95.55 91.01 90.61 96.56 88.08
@@ 4.2 Performance in Conflict-Resistance
Conflict-resistance regarding task number and difficulty. To further compare and analyze the conflict-resistance capabilities of different methods, we conduct comparative experiments on CSR-MTL by varying the number and difficulty of the tasks. In particular, for each task in CSR-MTL, we first calculate the average of all the methods’ performance based on the results in Table 5. The average performance of the methods in a task measures the difficulty of the task for the base model — the lower the average performance is, the more difficult the task is. Then, we sort the tasks in ascending order based on their difficulty. Finally, we adapt the base model for the top-$K$ tasks, K=2,,9K=2,...,9, and show the performance of different methods in Figure 1(b). With the increase of task number and difficulty, all the methods suffer performance degradation because $i)$ task conflict becomes severe as the number of tasks increases, and ii)ii) difficult tasks are more likely to have conflicts with other tasks. Notably, MoORE outperforms all other baselines across all settings.
(a) MMLU
(b) IFEval
(c) BBH
(d) GPQA
(e) HumanEval
(f) MBPP
(g) GSM-8K
(h) Overall degradation
Figure 2: The loss of performance in OR-MTL. (a-g) The gray dashed line represents the performance of LLaMA-3.1 8B before adaptation. (h) The overall performance degradation across all tasks.
@@ 4.3 Performance in Oblivion-Resistance
To demonstrate the superiority of MoORE in mitigating task oblivion, we compare various methods in the Language Model Evaluation Harness framework eval-harness . In particular, we first adapt the base model on CSR-MTL using different methods. Then, we evaluate the adapted models on OR-MTL, comparing them with the base model. Figure 2 shows that MoORE consistently mitigates task oblivion across all seven tasks of OR-MTL, with an average performance drop of only 1.31% compared to the base model. It significantly outperforms the other adaptation methods. On HumanEval, MoORE achieves performance exceeding the original model, with an improvement of 9.75%. Similarly, LoRAMoE and MTL-LoRA also obtain slight improvements of 4.27% and 0.16%, respectively. This intriguing phenomenon may imply that the datasets in CSR-MTL have some relevance to HumanEval, providing information and capabilities beneficial for solving this task.
Oblivion-resistance regarding task number and difficulty. Furthermore, we conduct experiments to investigate how the ability of oblivion-resistance changes with increased task number and difficulty. The results in Figure 1(c) show that after adapting to the top-3 tasks of CSR-MTL, MoORE and its strongest competitor, LoRAMoE dou2024loramoe , achieve comparable performance on OR-MTL. However, with the increase of task number and difficulty, the performance of LoRAMoE changes slightly while the performance of MoORE increases and becomes superior to that of LoRAMoE. This result demonstrates that MoORE has stronger oblivion-resistance than its competitors.
@@
The generalization to unseen task IDs. The computation of task-level weights requires task IDs as input, but the ID of a new task is unavailable during inference. Therefore, we use the average of the existing task ID embeddings as the embedding of the new task. Specifically, after adapting the base model on CSR-MTL, MoORE only learns the task IDs of the nine tasks in CSR-MTL. During evaluation on OR-MTL, we use the average of the nine task ID embeddings as the task ID embedding of each new task. The results in Figure 1(c) and 2 validate the effectiveness of this approach.
In addition, this experiment also explains the result in Figure 1(c). In particular, the more tasks considered in the adaptation phase, the more likely some tasks correlate with those covered in the pre-training phase. As a result, increasing the number of tasks in the adaptation phase helps enhance the oblivion-resistance of MoORE.
@@ 4.4 Inference Efficiency
To compare inference efficiency, we adapt the base model using MoORE and other baselines, and evaluate their inference time on GPQA. For a fair comparison, all experiments are conducted on a single NVIDIA A100 GPU with the same batch size. The results in Figure 1(d) show the average inference time per batch for each method, which indicates that MoORE achieves superior inference efficiency compared to most baselines and is comparable to MixLoRA. Compared to the base model (i.e., the gray dashed line), MoORE moderately increases inference time.
@@ 5 Conclusion and Future Work
Limitations and future work. As shown in Figure 1(d), the inference efficiency of MoORE is comparable to that of baselines. However, further reducing its complexity for large-scale applications is challenging. In particular, MoORE applies many more experts than classic MoE models, and the analytic experiments in Appendix C.4 show that it learns dense weights for the experts. In addition, besides imposing orthogonal adapters, we need to find a more effective solution to enhance model capacity in broader scenarios. In theory, the results in Figures 1(b) and 1(c) show that increasing the number of tasks leads to a higher risk of task conflict while helping mitigate task oblivion. How to understand this “trade-off”? How can the limitation in multi-task adaptation be quantified regarding the number of tasks? These are still open problems. Based on the above analysis, we plan to explore efficient implementations of MoORE and study its theoretical properties in the future.